Use completing the square to rewrite this equation in vertex form. x^2 +10x=8
Answer by ewatrrr(10682)

 
Hi,
Use completing the square to rewrite this equation in vertex form.
x^2 +10x=8
y = x^2 +10x- 8
y = (x+5)^2 - 25 - 8
y = (x+5)^2 - 33